# An interpretation of the Brownian bridge as a physics-informed prior for the Poisson equation

### v1

@@ 4.2 Finite-dimensional representations
Let $d>1$, $(\psi_{i})_{i\in\mathbb{N}}$ be an orthonormal basis for $L^{2}(\mathbb{R}^{d})$, and for each $n\in\mathbb{N}$, let $\mu^{(n)}$ be given by eq. (4.2). Then there exists a Lévy white noise $\mu$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ such that $\mu^{(n)}$ converges in distribution to $\mu$. $\mu$ has characteristic function $\phi_{\mu}(u)=\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}$, $u\in\mathcal{S}$.
@@
The proof can be found in Appendix B. ∎
Remark 4.1.
The space $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is inconveniently large in a scientific machine learning context. For instance, the Dirac delta distribution belongs to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, and we cannot make sense of generating such a sample. However, there are often smaller function spaces with full measure under a Lévy white noise which are easier to characterize. A general methodology for identifying a Besov space where this property holds is provided in [5]. In this case, the distribution is simply the $d$-dimensional Brownian sheet, conditioned to be zero on the boundaries. One can show this process is continuous on the unit cube [1].
@@
Theorem 4.2 and Theorem 4.4 justify the use of eq. (4.2) in applications. In 1D, the finite-dimensional representation of the prior converges to the correct Gaussian measure on $L^{2}(\Omega)$. We can use this form to derive additional results in the next section. In the multidimensional case, eq. (4.2) converges to the correct stochastic process. It is important to note that this cannot be taken as a Gaussian measure, as the covariance operator is not trace-class. However, we can still generate samples which approximate the prior using this finite-dimensional representation.
@@ 4.3 Convergence properties
We now discuss conditions for which the posterior converges to the ground truth and in what sense. Thankfully, understanding the convergence behavior is fairly straightforward due to the work of [45]. By applying the theorems derived in that work, we can prove that the posterior mean function will converge to the ground truth in the limit of infinite observations. This holds even if we estimate the hyperparameters of the prior. This fact is very relevant for us, since the source term $q$ could be treated as an unknown hyperparameter to the physics-informed prior. Again, we will start with the case $d=1$ to illustrate.
To begin, we must discuss a bit about how the data should be collected in order for the convergence conditions to hold. Specifically, we must characterize how uniformly the data points are collected in the domain. Let the set $X_{n}=(x_{1},\dots,x_{n})\subset\Omega$ represent the points at which the measurements are collected. The fill distance is defined by
@@
Both $h_{X_{n}}$ and $r_{X_{n}}$ go to 00 as $n\to\infty$ under a space-filling design, for example the uniform grid or Sobol net [37, 47]. In what follows, we will assume the measurements are collected on a uniform grid, so that $\rho_{X_{n}}$ is constant with $n$ and the calculations are simplified. The theorems also hold for any data collection scheme where $\rho_{X_{n}}$ is bounded above.
We rely on the two main convergence theorems from [45]. Notably, the results are concerned with the case where the GP prior contains unknown hyperparameters which are approximated along with the field. In that work, empirical Bayes is taken as the motivating example. The conditions on the hyperparameters are fairly loose and the convergence theorems will hold in a wide variety of cases. The way in which the hyperparameters are learned does not impact the results, and one may prefer an alternative such as a maximum likelihood estimate (MLE). We put a specific focus on the MAP estimate in Section 5.
@@
Theorem 4.5 (Theorem 3.5 [45]).
@@
  3. (iii)
    $u^{*}\in H^{\bar{\tau}}(\Omega)$, for some $\bar{\tau}=\alpha+\gamma$ with $\alpha\in\mathbb{N}$, $\alpha>d/2$, and $10\leq\gamma<$ 1.
  4. (iv)
@@
  5. (v)
    For some $N^{*}\in\mathbb{N}$, the quantities $\tau^{-}=\inf_{n\geq N^{*}}\tau(\hat{\lambda}_{n})$ and $\tau^{+}=\sup_{n\geq N^{*}}\tau(\hat{\lambda}_{n})$ satisfy $\tilde{\tau}=\alpha^{\prime}+\gamma^{\prime}$ with $\alpha^{\prime}\in\mathbb{N}$, $\alpha^{\prime}>d/2$ and $10\leq\gamma^{\prime}<$ 1.
@@
We discuss some of the assumptions of Theorem 4.5 in the context of our problem. The third assumption is a regularity constraint on the ground truth. Since we are mostly concerned with identifying the solution to the Poisson equation, this is reasonable to impose. Assuming that $u^{*}$ is a solution to the Poisson equation (for sufficiently regular domain and source term), we would expect at the minimum $u^{*}\in H^{2}(\Omega)$, which satisfies (iii) up to $d=3$, e.g. picking $\gamma=0.5$ when $d=3$. Observe that we may have convergence for any sufficiently smooth ground truth field, not just solutions to the assumed PDE. This is relevant, for instance, in the case of model-misspecification. This could result from an incorrectly identified source or perhaps $u^{*}$ is better modeled by the nonlinear Poisson equation, among others.
Assumption (iv) is a regularity constraint on the prior mean function, $u^{0}$. As with assumption (iii), this is easy to satisfy, as $u^{0}(\cdot;\lambda)$ is exactly a solution to the Poisson equation for any $\lambda$. As an example, if we represent $q$ as a neural network with a smooth activation function, then this assumption trivially holds, even as the network weights are updated.
The final assumption is related to how the hyperparameters are learned. The quantities $\tau^{-}$ and $\tau^{+}$ are essentially $\lim\inf\tau(\hat{\lambda}_{n})$ and $\lim\sup\tau(\hat{\lambda}_{n})$. This assumption simply requires the the RKHS of the prior covariance to remain sufficiently smooth as the hyperparameters are optimized, and immediately holds if $\lambda$ is kept fixed. In our physics-informed prior, $q$ does not enter the covariance, so this is only an assumption on $\beta$. Restricting $\beta$ to $0<\beta<\infty$ will satisfy this condition, as the RKHS does not change as $\beta$ moves. We encourage the reader to refer to [45] for details on optimal convergence rates.
@@
Let $\Omega=[0,1]$, $q(\cdot;\theta)\in L^{2}(\Omega)$ for all $\theta$, $u^{*}\in H^{2}(\Omega)$, and $u^{0}(\cdot;\theta)$ be the solution to (1.1). Take $\hat{\lambda}\subset\Lambda$ to be a sequence of estimates for the collection $(\theta,\beta)$ for compact $\Lambda\subseteq\mathrm{dom}(\lambda)$. Then the GP posterior mean function, $\tilde{m}(\cdot;\hat{\lambda}_{n})$, given by eq. (2.4), with prior $u(\cdot;\hat{\lambda}_{n})\sim\mathcal{GP}(u^{0}(\cdot;\hat{\theta}_{n}),(-\hat{\beta}_{n}\Delta)^{-1})$, converges in $L^{2}(\Omega)$ to $u^{*}$ in the limit of infinite observations. That is,
@@
An immediate corollary is the following, which may be of interest in physics-informed machine learning tasks.
@@
Let $E(u;\theta)=\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}+q(\cdot;\theta)u\>d\Omega$, $\hat{\eta}_{n}=\sigma^{2}\hat{\beta_{n}}/n$, and the assumptions of Theorem 4.6 hold. Then,
@@
where $\hat{u}(\cdot,\hat{\lambda}_{n})$ is the solution to the physics-informed optimization problem
$\hat{u}(\cdot;\hat{\lambda}_{n})=\operatorname*{arg\,min}_{u\in H_{k}}\mathcal{L}_{\mathrm{data}}(u)+\hat{\eta}_{n}E(u;\hat{\theta}_{n}).$
@@
The hypotheses of Theorem 4.5 are the exactly the same as what is found in [45, Theorem 3.8], which shows there exists a constant $c$, independent of $n$, with
@@
The above results are valid for the case $d=1$. Similar convergence theorems also hold in the general case, but one must instead rely on [45, Theorem 3.11], which exploits the tensor product structure of the covariance kernel. Note that in order for the results to hold for $d>1$, a sparse grid data collection scheme must be used.
We now mention some implications of Theorem 4.6. The first observation is that convergence holds even under significant model-form error. In practice we a priori assume the ground truth satisfies the Poisson equation. If we have selected the wrong model, i.e. the Poisson equation does not model the system accurately, then convergence still holds provided that the ground truth satisfies some smoothness constraints. The same is true for model-form error resulting from picking the wrong source term or incorrectly identifying $q$ if we are solving the inverse problem.
The assumptions on $q$ and $\beta$ are rather loose in the application of this theorem. When solving the inverse problem, the conditions on Theorem 4.6 may be satisfied even if we have identified a bad estimate for $q$. In fact, $q$ need not be identifiable. The main requirement is that $\lambda$ remains in a compact domain. If we represent $q$ with a neural network, this is satisfied if we do not allow the weights to explode. Unfortunately, there is no guarantee that our estimate of $q$ (or $\beta$) will converge to the truth. We leave the discussion on this to Section 5.

### v2

@@ 4.2 Finite-dimensional representations
Let $d=1$, $(\psi_{i})_{i\in\mathbb{N}}$ be an orthonormal basis of $L^{2}(\Omega)$, and $\mu_{C}=\mathcal{N}(0,C)$ with corresponding posterior $\mu_{C}^{y}$ as given by Theorem 3.3. For each $n\in\mathbb{N}$, let $\mu_{n}$ be given by eq. (4.2) with posterior $\mu_{n}^{y}$ eq, (4.3). Further, assume that $\Phi:L^{2}(\Omega)\to[0,\infty)$ is continuous and measurable. Then for any $p\geq 1$, we have $W_{p}(\mu_{C}^{y},\mu_{n}^{y})\to 0$ as $n\to\infty$.
@@
The proof is a simple application of [sprungk2020local, Lemma 16], which states that $W_{p}$ convergence of the priors implies convergence in the posteriors. Combining Lemma 4.2 with [villani2008optimal, Theorem 6.9], we have $W_{p}(\mu_{C},\mu_{n})\to 0$, which gives the result. ∎
Next, we move to the general case $d>1$. Recall that in this setting the Brownian bridge does not define a Gaussian measure on $L^{2}(\Omega)$ due to the fact that the associated covariance is not a trace-class operator under the $L^{2}$-norm. However, we can still provide parallel convergence theorems as with the case $d=1$ if we adjust the function spaces involved. One option is to work in the setting of tempered distributions, which is discussed in Appendix C. This is common in white noise analysis [hida2013white], e.g. stochastic PDEs [da2014stochastic], and mathematical physics [glimm2012quantum]. Instead, we will define the prior on the dual of an appropriate Sobolev space, otherwise known as a negative Sobolev space.
Recall that for a $\tau\in\mathbb{R}$, the dual of $H^{\tau}(\Omega)$ is given by
$\mathcal{H}^{-\tau}(\Omega)=\left\{u:H^{\tau}(\Omega)\to\mathbb{R}:\|u\|^{2}_{H^{-\tau}}\coloneqq\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{\tau}\langle u,\psi_{\alpha}\rangle^{2}<\infty\right\},$
where again $\phi_{\alpha}$ and $\lambda_{\alpha}$ are the eigenfunctions with corresponding eigenvalues of the Brownian bridge, as given by eq. (3.2) and eq. (3.3), respectively [adams2003sobolev]. First, we show that the covariance of the Brownian bridge has finite trace on $H^{-\tau}(\Omega)$ when $\tau>d/2-1$.
Proposition 4.3.
Let $C$ be the covariance operator of the Brownian bridge with $d>1$. Then $C$ defines a trace class operator on $H^{-\tau}(\Omega)$ if and only if $\tau>d/2-1$.
Proof.
On the space $H^{-\tau}(\Omega)$, the trace class norm is $\|C\|_{\mathrm{tr}\left(H^{-\tau}\right)}=\sum_{|\alpha|\in\mathbb{N}}\langle C\psi_{\alpha},\psi_{\alpha}\rangle_{H^{-\tau}}.$ Choosing the basis to be the eigenfunctions of $C$, an individual entry of the series is
$c_{\alpha}$ $=\langle C\psi_{\alpha},\psi_{\alpha}\rangle_{H^{-\tau}}$
$=\langle\lambda_{\alpha}\psi_{\alpha},\psi_{\alpha}\rangle_{H^{-\tau}}$
$=\lambda_{\alpha}\|\psi_{\alpha}\|^{2}_{H^{-\tau}}$
$=\lambda_{\alpha}^{1+\tau}\langle\psi_{\alpha},\psi_{\alpha}\rangle^{2}=\lambda_{\alpha}^{1+\tau}.$
Plugging in the form of the eigenvalues, we have $\|C\|_{\mathrm{tr}\left(H^{-\tau}\right)}=\sum_{|\alpha|\in\mathbb{N}}(\pi|\alpha|)^{-2(1+\tau)}$, which converges if and only if $2(1+\tau)>d$ by comparison to an integral. ∎
Proposition 4.3 shows that the physics-informed prior is a well-defined Gaussian measure on $H^{-\tau}(\Omega)$ for $\tau$ sufficiently large, and we do not need to restrict ourselves to the setting of GPs. This means that we are justified in writing $\mu_{C}\sim\mathcal{N}(0,C)$. Unfortunately, these spaces are a bit larger than we might like, given that $H^{-\tau}(\Omega)$ contains distributions. We also have the inclusion relation $H^{\tau}(\Omega)\subset L^{2}(\Omega)\subset H^{-\tau}(\Omega)$. In a machine learning context, it is difficult to make sense of generating such a sample. Fortunately, the Brownian bridge is known to produce continuous sample paths on the unit cube [adler2010geometry].
There are certain advantages gained by working in this space, however. Namely that, since the prior is indeed a probability measure, we gain access to both Bayes’s theorem and the Wasserstein distance. In order to apply Bayes’s theorem, we must ensure that the domain of the observation map $R$ is a space with full measure under $\mu_{C}$. This holds when $R$ takes continuous functions as input, which we have assumed in view of Proposition 3.1. Hence, by Bayes’s rule we identify a posterior measure $\mu^{y}$ and in similar fashion a finite-dimensional approximation $\mu^{y}_{n}$, both viewed as probability measures over $H^{-\tau}(\Omega)$. With that out of the way, we provide convergence criteria, starting with the prior. The first step is to show convergence of the covariance forms.
Lemma 4.3.
Let $d>1$, $\tau>d/2-1$, $(\psi_{\alpha})_{|\alpha|\in\mathbb{N}}$ be the eigenfunctions of $C$, and for each $n$, $P_{n}$ be the projection of $H^{-\tau}(\Omega)$ onto the finite span $\{\psi_{\alpha}\}_{|\alpha|=1}^{n}$, i.e. $\mathcal{F}_{n}=P_{n}(H^{-\tau}(\Omega))$. Then the finite rank covariances $\Sigma_{\mathcal{F}_{n}}=P_{n}CP_{n}$ satisfy $\|C-\Sigma_{\mathcal{F}_{n}}\|_{\mathrm{tr}\left(H^{-\tau}\right)}\to 0$.
Proof.
Beyond the order of the projection, $n$, the operators $C$ and $C-\Sigma_{\mathcal{F}_{n}}$ have the same eigenvalues. So, $\|C-\Sigma_{\mathcal{F}_{n}}\|_{\mathrm{tr}\left(H^{-\tau}\right)}=\sum_{|\alpha|>n}(\pi|\alpha|)^{-2(1+\tau)}$, which is just the tail of the same convergent series from the proof of Proposition 4.3. ∎
This shows that by projection onto the eigenfunctions, the finite-dimensional approximation we have chosen can also be viewed as a Gaussian measure on $H^{-\tau}$. Noting that Gaussian measures are uniquely determined by their covariance forms, Lemma 4.3 also shows that $\mu_{n}$ converges in some sense to the correct measure. In the next result, we make this precise in terms of $W_{2}$ distance.
Theorem 4.4.
Let $d>1$, $\tau>d/2-1$, and for each $n\in\mathbb{N}$, $P_{n}$ be the projection onto the finite span of eigenfunctions $\{\psi_{\alpha}\}_{|\alpha|=1}^{n}$. Further, let $\mu_{n}$ be given by eq. (4.2) with $\mathcal{F}_{n}=P_{n}(H^{-\tau}(\Omega))$ with corresponding posterior $\mu^{y}_{n}$ following eq. (4.3). Then in the limit $n\to\infty$, it holds that both $W_{2}(\mu_{C},\mu_{n})\to 0$ and $W_{2}(\mu^{y},\mu^{y}_{n})\to 0$.
Proof.
The result follows from application of Gelbrich’s formula to the priors
$W^{2}_{2}(\mu_{C},\mu_{n})=\|C\|_{\mathrm{tr}\left(H^{-\tau}\right)}+\|\Sigma_{\mathcal{F}_{n}}\|_{\mathrm{tr}\left(H^{-\tau}\right)}-2\left\|\sqrt{C^{1/2}\Sigma_{\mathcal{F}_{n}}C^{1/2}}\right\|_{\mathrm{tr}\left(H^{-\tau}\right)}.$
It holds from continuity of the map $A\mapsto A^{1/2}$, continuity of the norm, and by Lemma 4.3 that $\|\Sigma_{\mathcal{F}_{n}}\|_{\mathrm{tr}\left(H^{-\tau}\right)}\to\|C\|_{\mathrm{tr}\left(H^{-\tau}\right)}$ and
$\left\|\sqrt{C^{1/2}\Sigma_{\mathcal{F}_{n}}C^{1/2}}\right\|_{\mathrm{tr}\left(H^{-\tau}\right)}\to\|C\|_{\mathrm{tr}\left(H^{-\tau}\right)},$
which can be checked by spectral representation of the operators. Therefore in the limit $W_{2}(\mu_{C},\mu_{n})\to 0$. Posterior convergence follows immediately from [sprungk2020local, Lemma 16]. ∎
Of course, an immediate corollary is that $\mu_{n}^{y}\implies\mu^{y}$ by [villani2008optimal, Theorem 6.9].
@@
Theorems 4.3 and 4.4 justify the use of eq. (4.2) in applications. In 1D, the finite-dimensional representation of the prior converges to the correct Gaussian measure on $L^{2}(\Omega)$. We can use this form to derive additional results in the next section. In the multidimensional case, eq. (4.2) converges to the correct measure on the dual of a Sobolev space.
@@ 4.3 Convergence properties
We now discuss conditions for which the posterior converges to the ground truth and in what sense. Thankfully, understanding the convergence behavior is fairly straightforward due to the work of [teckentrup2020convergence]. By applying the theorems derived in that work, we can prove that the posterior mean function will converge to the ground truth in the limit of infinite observations. This holds even if we estimate the hyperparameters of the prior. This fact is very relevant for us, since the source term $q$ could be treated as an unknown hyperparameter to the physics-informed prior. Again, we will start with the case $d=1$ to illustrate.
To begin, we must discuss a bit about how the data should be collected in order for the convergence conditions to hold. First, we restrict ourselves to point measurements in the domain $\Omega$. Then, we must characterize how uniformly the data points are collected in the domain. Let the set $X_{n}=(x_{1},\dots,x_{n})\subset\Omega$ represent the points at which the measurements are collected. The fill distance is defined by
@@
Both $h_{X_{n}}$ and $r_{X_{n}}$ go to 0 as $n\to\infty$ under a space-filling design, for example the uniform grid or Sobol net [niederreiter1992random, wynne2021convergence]. In what follows, we will assume the measurements are collected on a uniform grid, so that $\rho_{X_{n}}$ is constant with $n$ and the calculations are simplified. The theorems also hold for any data collection scheme where $\rho_{X_{n}}$ is bounded above.
We rely on the two main convergence theorems from [teckentrup2020convergence]. Notably, the results are concerned with the case where the GP prior contains unknown hyperparameters which are approximated along with the field. In that work, empirical Bayes is taken as the motivating example. The conditions on the hyperparameters are fairly loose and the convergence theorems will hold in a wide variety of cases. The way in which the hyperparameters are learned does not impact the results, and one may prefer an alternative such as a maximum likelihood estimate (MLE). We put a specific focus on the MAP estimate/evidence approximation in Section 5.
@@
Theorem 4.5 (Theorem 3.5 [teckentrup2020convergence]).
@@
  3. 3.
    $u^{*}\in H^{\bar{\tau}}(\Omega)$, for some $\bar{\tau}=\alpha+\gamma$ with $\alpha\in\mathbb{N}$, $\alpha>d/2$, and $0\leq\gamma<1$.
  4. 4.
@@
  5. 5.
    For some $N^{*}\in\mathbb{N}$, the quantities $\tau^{-}=\inf_{n\geq N^{*}}\tau(\hat{\lambda}_{n})$ and $\tau^{+}=\sup_{n\geq N^{*}}\tau(\hat{\lambda}_{n})$ satisfy $\tilde{\tau}=\alpha^{\prime}+\gamma^{\prime}$ with $\alpha^{\prime}\in\mathbb{N}$, $\alpha^{\prime}>d/2$ and $0\leq\gamma^{\prime}<1$.
@@
We discuss some of the assumptions of Theorem 4.5 in the context of our problem. The third assumption is a regularity constraint on the ground truth. Since we are mostly concerned with identifying the solution to the Poisson equation, this is reasonable to impose. Assuming that $u^{*}$ is a solution to the Poisson equation (for sufficiently regular domain and source term), we would expect at the minimum $u^{*}\in H^{2}(\Omega)$ in light of Proposition 3.1, which satisfies (iii) up to $d=3$, e.g. picking $\gamma=0.5$ when $d=3$. Observe that we may have convergence for any sufficiently smooth ground truth field, not just solutions to the assumed PDE. This is relevant, for instance, in the case of model-misspecification. This could result from an incorrectly identified source or perhaps $u^{*}$ is better modeled by the nonlinear Poisson equation, among others.
Assumption (iv) is a regularity constraint on the prior mean function, $u^{0}$. As with assumption (iii), this is easy to satisfy, as $u^{0}(\cdot;\lambda)$ is exactly a solution to the Poisson equation for any $\lambda$. As an example, if we represent $q$ as a neural network with a smooth activation function or as a GP with smooth sample paths, then this assumption trivially holds, even as the network weights are updated.
The final assumption is related to how the hyperparameters are learned. The quantities $\tau^{-}$ and $\tau^{+}$ are essentially $\lim\inf\tau(\hat{\lambda}_{n})$ and $\lim\sup\tau(\hat{\lambda}_{n})$. This assumption simply requires the the RKHS of the prior covariance to remain sufficiently smooth as the hyperparameters are optimized, and immediately holds if $\lambda$ is kept fixed. In our physics-informed prior, $q$ does not enter the covariance, so this is only an assumption on $\beta$. Restricting $\beta$ to a compact interval will satisfy this condition, as the RKHS does not change as $\beta$ moves. We encourage the reader to refer to [teckentrup2020convergence] for details on optimal convergence rates.
@@
Let $\Omega=[0,1]$, $q(\cdot;\theta)\in L^{2}(\Omega)$ for all $\theta$, $u^{*}\in H^{2}(\Omega)$, and $u^{0}(\cdot;\theta)$ be the solution to eq. (1.1). Take $\hat{\lambda}_{n}\subset\Lambda$ to be a sequence of estimates for the collection $(\theta,\beta)$ for compact $\Lambda\subseteq\mathrm{dom}(\lambda)$. Then the GP posterior mean function, $\tilde{m}(\cdot;\hat{\lambda}_{n})$, given by eq. (2.4), with prior $u(\cdot;\hat{\lambda}_{n})\sim\mathcal{GP}(u^{0}(\cdot;\hat{\theta}_{n}),(-\hat{\beta}_{n}\Delta)^{-1})$, converges in $L^{2}(\Omega)$ to $u^{*}$ in the limit of infinite observations. That is,
@@
An immediate corollary is the following, which we state in terms of the variational problem we started with.
@@
Let $E(u;\theta)=\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}+q(\cdot;\theta)u\>d\Omega$, $\hat{r}_{n}=\sigma^{2}\hat{\beta_{n}}/n$, and the assumptions of Theorem 4.6 hold. Then,
@@
where $\hat{u}(\cdot,\hat{\lambda}_{n})$ is the solution to the physics-regularized inverse problem
$\hat{u}(\cdot;\hat{\lambda}_{n})=\operatorname*{arg\,min}_{u\in H_{k}}\mathcal{L}_{\mathrm{data}}(u)+\hat{r}_{n}E(u;\hat{\theta}_{n}).$
@@
The hypotheses of Theorem 4.5 are the exactly the same as what is found in [teckentrup2020convergence, Theorem 3.8], which shows there exists a constant $c$, independent of $n$, with
@@
The above results are valid for the case $d=1$. Similar convergence theorems also hold for $d>1$, but one must instead rely on [teckentrup2020convergence, Theorem 3.11], which exploits the tensor product structure of the covariance kernel. Note that in order for the results to hold for $d>1$, a sparse grid data collection scheme must be used.
We now mention some implications of Theorem 4.6. The first observation is that convergence holds even under significant model-form error. In practice we a priori assume the ground truth satisfies the Poisson equation. If we have selected the wrong model, i.e., the Poisson equation does not model the system accurately, then convergence still holds provided that the ground truth satisfies some smoothness constraints. The same is true for model-form error resulting from picking the wrong source term or incorrectly identifying $q$ if we are solving the inverse problem.
The assumptions on $q$ and $\beta$ are rather loose in the application of this theorem. When solving the inverse problem, the conditions on Theorem 4.6 may be satisfied even if we have identified a bad estimate for $q$. In fact, $q$ need not be identifiable. The main requirement is that $\lambda$ remains in a compact domain. If we represent $q$ with a neural network, this is satisfied if we do not allow the weights to explode. Unfortunately, we are unable to prove if an estimate of $q$ will also converge to the ground truth. We leave the discussion on this to Section 5.